# Normality-Guided Distributional Reinforcement Learning for Continuous Control

### v3

@@ 4.1 State Distributional Value Function
We first similarly introduce a state distributional value function $V^{D,\pi}(s)$ that also outputs quantile bars for state $s$. The distributional value function $V^{D,\pi}(s)$ for the general actor-critic policy gradient is only dependent on the state. To be specific, the target for $V^{D,\pi}(s_{t})$ is $R(s_{t},a_{t})+V^{D,\pi}(s_{t+1})$ similar to Equation (4). The quantile bars of $V^{D,\pi}(s_{t})$ is shifted by the amount of $R(s_{t},a_{t})$. This let us use the actor-critic algorithms such as Policy Optimization (PPO) Schulman et al. (2017) and Trust Region Policy Optimization (TRPO) Schulman et al. (2015a).
@@ 4.2 Analysis of Distributional Value Function
Figure 3: The standard deviation computed from each of the distributional value functions trained on the distributional Bellman optimality operator (Equation 7) with exponential smoothing. The estimated standard deviation increases as timestep increases.
Furthermore, our observation reveals that the learned $V^{D,\pi}$ and $Q^{D,\pi}$ exhibit an increase in variances as timesteps progress, contradicting Theorem 1. The distributional Bellman optimality operator Bellemare et al. (2017b) causes the variance of the distributional value function to increase with timestep as the discount factor shrinks quantiles. To illustrate, suppose $Q^{D,\pi}(s_{t+1},a_{t+1})$ = $\{q_{0},q_{1},...,q_{N-1}\}$ at timestep ${t+1}$. Then, the target for $Q^{D,\pi}(s_{t},a_{t})$ under the distributional Bellman operator will be $\{R(s_{t},a_{t})+\gamma q_{0},R(s_{t},a_{t})+\gamma q_{1},...,R(s_{t},a_{t})+\gamma q_{N-1}\}$, i.e., $\{q_{0},q_{1},...,q_{N-1}\}$ is shifted by $R(s_{t},a_{t})$ and shrunk by $\gamma<1$. Consequently the range of the distribution for timestep $t$ is narrower than the range for timestep $t+1$.
We test this empirically by training distributional value functions with PPO and QR-DQN using the distributional Bellman optimality operator (Equation 7) and graphing the estimated standard deviation from quantile output for each state (Figure 3). Indeed, the variance estimates appear to increase with timestep on average. We now propose a method that utilizes the variance network and returns to generate quantiles for the distributional value function.
@@ 4.3 Variance of Normal Distribution
Variance network ($\sigma^{2}_{\psi}$), introduced by Nix and Weigend Nix and Weigend (1994) and named by Mai et al. Mai et al. (2022), are designed to capture both the mean and Gaussian error (variance). In line with the normal assumption, the variance network allows us to estimate the variance of a Gaussian distribution. We leverage a variance network to estimate a variance for each state when computing quantile bars.
@@
$\mathcal{L}_{\sigma^{2}}(\psi)=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{2\sigma^{2}_{\psi}(s_{i})}||q_{\mathrm{{avg}}}(s_{i})-G(s_{i})||^{2}+\frac{1}{2}\rm{ln}\sigma^{2}_{\psi}(s_{i})$
where $q_{\mathrm{{avg}}}(s)$ is the mean of $V^{D,\pi}_{\theta}(s)$ and $G(s)$ is the sum of sampled rewards from the state $s$ (return). This separation of prediction eliminates the need for additional hyperparameter tuning that would have been required for the simultaneous prediction of mean and variance, as observed in Mai et al. Mai et al. (2022). Although the numerator of Equation 9 is separated from $\psi$, the update is performed in the direction of increasing $\sigma^{2}_{\psi}$ when the error of the numerator is significant. Conversely, if the error is sufficiently small, $\sigma^{2}_{\psi}$ is decreased. $\sigma^{2}_{\psi}$ is trained to predict variance in a manner consistent with the original variance network.
@@ 4.4 Target for Distributional Value Function
We provide the quantile bars of an approximated normal distribution to $V^{D,\pi}_{\theta}$ with the variance from $\sigma^{2}_{\psi}$ above and returns. To compute the target quantile bars, we need the mean of the approximated normal distribution. The mean can be TD target or the sum of rewards $G(s)$ from $s$ to end. Here, we use return $G$ for the sake of notation simplicity. To obtain a quantile target for $V^{D,\pi}$, we first precompute the $Z$ values of the standard normal distribution. For example, when $N=4$, $Z=\{-0.841,-0.253,0.253,0.841\}$, i.e., $P(X<=Z[i])\approx 0.2*(i+1)=\frac{i+1}{N+1}$ for $i\in\{0,1,2,3\}$. Then we compute the target quantile bars $q^{\prime}_{t}$ of $V^{D,\pi}_{\theta}(s_{t})$ such that $q^{\prime}_{t}=\{R_{t}+\sigma_{\psi}Z[i]\,|\,i\in\{0,1,...,N-1\}\}$ and fit $V^{D,\pi}_{\theta}$ to $q^{\prime}_{t}$ by quantile Huber loss Aravkin et al. (2014). The quantile Huber loss makes $V^{D,\pi}_{\theta}$ output quantile bars that match the definition of quantiles. The Huber loss Huber (1964) is as follows:
@@ 4.5 Uncertainty Weight
1:  Input: policy $\pi$, distributional value function $V^{D,\pi}_{\theta}$,
2:             variance network $\sigma^{2}_{\psi}$, rollout buffer $B$
3:  Initialize $\pi$, $V^{D,\pi}$, and $\sigma^{2}_{\psi}$
4:  for $i=1$ to epoch_num
5:     for $j=1$ to rollout_num
6:        $a_{t}\sim\pi(\cdot|s_{t})$
7:        $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$
8:        Compute $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$ with
9:            $V^{D,\pi}(s_{t})=\{q_{0}(s_{t}),q_{1}(s_{t}),...,q_{N-1}(s_{t})\}$
10:        Find $q^{\prime}(s_{t})$ from $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$
11:        Compute the mean absolute error
12:           $E=\sum_{i=0}^{N-1}|q_{i}(s_{t})-q_{i}^{\prime}(s_{t}))|$
13:        Store ($s_{t},a_{t},r(s_{t},a_{t}),E,V^{D,\pi}(s_{t}),\sigma^{2}_{\psi}(s_{t})$) in $B$
14:        if env is done
15:           Reset env
16:     end for
17:     Perform the parametric search to find
18:        the temperature $T$ (Appendix B)
19:     Compute target quantiles for $V^{D,\pi}$
20:        with the stored return and $\sigma^{2}_{\psi}$
21:     Minimize $\mathcal{L}_{\sigma^{2}}(\psi)$ and $\mathcal{L}_{q}(\theta)$
22:     Optimize $\pi$ with a policy objective scaled by $w$
23:  end for
Table 1: Ablation study of TRPO and PPO. Each entry represents an average return and a standard error from over 30 episodes with 30 different settings. $V^{D,\pi}$ (QR) represents the distributional value function trained with the Huber quantile regression. This table illustrates an improvement trend across the scores with the addition of each component from the baseline to MC-CLT. Also, MC-CLT results generally better than the Ensemble method, while using fewer weights and reduced training time. An asterisk (*) indicates a statistically significant improvement between MC-CLT and Ensemble method with a $p$-value of $0.05$.
@@
Swimmer Walker2d InvertedDoublePendulum LunarLanderContinuous
@@
Swimmer Walker2d InvertedDoublePendulum LunarLanderContinuous
@@
Figure 2 illustrates this when the number of quantile bars is $N=4$. If the prediction of $V^{D,\pi}$ is reliable, it should be symmetrical with respect to the mean of the quantile bars from our normality assumption, and the location of the quantile bars should follow a normal distribution. We first find the mean $q_{\mathrm{{avg}}}(s)$ = $\frac{1}{N}\sum_{i=0}^{i=N-1}q_{i}(s)$ of the quantile bars and calculate the standard deviations $\{\sigma_{0},\sigma_{1},...,\sigma_{N-1}\}$ for each quantile such that $P(Z\leq\frac{q_{i}-q_{\mathrm{{avg}}}}{\sigma_{i}})=\frac{i+1}{N+1}$ for $i\in\{0,1,...,N-1\}$. We consider $V^{D,\pi}$ is trying to approximate $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$ where $\sigma_{\mathrm{{avg}}}=\frac{1}{N}\sum_{i=0}^{i=N-1}\sigma_{i}$.
To measure the uncertainty $w$ for policy updates, we find the correct locations of each quantile $\{q_{0}^{\prime}(s),...,q_{N-1}^{\prime}(s)\}$ based on $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$. The better $V^{D,\pi}$ resembles the normal distribution, the smaller the difference between these two types of quantile bars. Therefore, we use the mean absolute error $E=\sum_{i=0}^{i=N-1}|q_{i}(s)-q_{i}^{\prime}(s)|$, and we compute the uncertainty weight $w(s)$ as follows:
$w(s)=\sigma(-E*T)+0.5$ (12)
where $T>0$ is a temperature and $\sigma$ is the sigmoid function. Although Lee et al. Lee et al. (2020) suggest values for $T$, it seems that finding the reasonable value for $T$ still requires some tuning. We instead set a target uncertainty weight and perform a parametric search by adjusting $T$ to achieve that target weight (Appendix B).
4.6 Policy Update with Uncertainty Weight
We leverage the uncertainty weights $w$ to update the policy to mainly focus on values for which the current $V^{D,\pi}$ predicts correctly. Each reinforcement learning algorithm has a policy objective function. We scale the objective function with the uncertainty weight to adjust the influence of the gradient for each sample. For TRPO Schulman et al. (2015a),
[…]
@@ 5.1 Experiment Setups
Our implementation is based on Spinning Up Achiam (2018). For our experiments, we chose two representative deep reinforcement learning algorithms, Policy Optimization (PPO) Schulman et al. (2017) and Trust Region Policy Optimization (TRPO) Schulman et al. (2015a), to evaluate our method. The advantage is computed by GAE Schulman et al. (2015b).
We use the default hyperparameters such as learning rate and batch size. All policies have a two-layer tanh network with 64 x 32 units, and all of the value function and the distributional value function has a two-layer ReLU network with 64 x 64 units or 128 x 128 units for all environments. All networks are updated with Adam optimizer Kingma and Ba (2014). We evaluate our method on continuous OpenAI gym Box2D Brockman et al. (2016) and MuJoCo tasks Todorov et al. (2012), as these environments have continuous action spaces and dense reward functions to use the normal approximation of MC-CLT (Theorem 1).
@@ 5.2 Detailed Ablation Analysis and Comparison
Table 1 represents the experimental results across 8 environments. The first section of the table provides results from policies trained using TRPO, while the second section gives results from policies trained with PPO. As we incrementally add each component from the baseline to MC-CLT, a consistent trend of performance improvement is observed. Also, our method shows better results compared to the ensemble-based approach in general. In 10 out of the 16 settings, we achieve statistically significant improvements between MC-CLT and the Ensemble method, as indicated by an asterisk (***) in the table, with a $p$-value of $0.05$. It’s worth noting that the ensemble-based approach with TRPO exhibits better mean performances in three settings: HalfCheetah, Ant, and InvertedDoublePendulum. However, the differences are not statistically significant.
@@
In addition, MC-CLT utilizes fewer weights and achieves faster training time, as shown in (Table 2). Table 2 presents a relative comparison with the other two methods based on the baseline. Specifically, when the training time for PPO is set to 1, the MC-CLT method requires 1.5 to 1.9 times more training time compared to the baseline. Similarly, the ensemble-based approach takes 2.8 to 3.1 times more time. Additionally, the MC-CLT method requires approximately 2.0 to 2.2 times more weights, while the ensemble-based approach uses 5 times as many parameters.
While MC-CLT requires hyperparameters for the number of quantile bars and the variance network, it necessitates fewer hyperparameters compared to previous methods Lee et al. (2020); Mai et al. (2022). We observe satisfactory performance when the number of quantiles is set to $\geq 8$ and the size of the variance network is equal to that of the distributional value function. Consequently, aside from the aspects controlling uncertainty, MC-CLT shares the same hyperparameters as the baseline method. The detailed hyperparameter settings are discussed in Appendix A.

### v4

@@ 4.1 State Distributional Value Function
We first similarly introduce a state distributional value function $V^{D,\pi}(s)$ that also outputs quantile bars for state $s$. The distributional value function $V^{D,\pi}(s)$ for the general actor-critic policy gradient is only dependent on the state. To be specific, the target for $V^{D,\pi}(s_{t})$ is $R(s_{t},a_{t})+V^{D,\pi}(s_{t+1})$ similar to Equation (4). The quantile bars of $V^{D,\pi}(s_{t})$ is shifted by the amount of $R(s_{t},a_{t})$. This let us use the actor-critic algorithms such as Policy Optimization (PPO) (Schulman et al., 2017) and Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a).
@@ 4.2 Analysis of Distributional Value Function
Furthermore, our observation reveals that the learned $V^{D,\pi}$ and $Q^{D,\pi}$ exhibit an increase in variances as timesteps progress, contradicting Theorem 1. The distributional Bellman optimality operator (Bellemare et al., 2017b) causes the variance of the distributional value function to increase with timestep as the discount factor shrinks quantiles. To illustrate, suppose $Q^{D,\pi}(s_{t+1},a_{t+1})$ = $\{q_{0},q_{1},...,q_{N-1}\}$ at timestep ${t+1}$. Then, the target for $Q^{D,\pi}(s_{t},a_{t})$ under the distributional Bellman operator will be $\{R(s_{t},a_{t})+\gamma q_{0},R(s_{t},a_{t})+\gamma q_{1},...,R(s_{t},a_{t})+\gamma q_{N-1}\}$, i.e., $\{q_{0},q_{1},...,q_{N-1}\}$ is shifted by $R(s_{t},a_{t})$ and shrunk by $\gamma<1$. Consequently the range of the distribution for timestep $t$ is narrower than the range for timestep $t+1$.
We test this empirically by training distributional value functions with PPO and QR-DQN using the distributional Bellman optimality operator (Equation LABEL:distributional_bellman_operator) and graphing the estimated standard deviation from quantile output for each state (Figure 2). Indeed, the variance estimates appear to increase with timestep on average. We now propose a method that utilizes the variance network and returns to generate quantiles for the distributional value function.
Figure 3: Illustration of how we measure uncertainty. When $N=4$, there are 4 quantile bars $\{q_{0},q_{1},q_{2},q_{3}\}$ as an output of $V^{D,\pi}$. Compute the mean $q_{\mathrm{{avg}}}$ of the quantile bars, and find $\sigma_{i}$ such that $P(Z\leq\frac{q_{i}-q_{\mathrm{{avg}}}}{\sigma_{i}})=0.2*(i+1)=\frac{i+1}{N+1}$ for $i\in\{0,1,2,3\}$. We consider $V^{D,\pi}$ as approximately $\mathcal{N}(q_{\mathrm{{avg}}},\sigma^{2}_{\mathrm{{avg}}})$ where $\sigma_{\mathrm{{avg}}}=\frac{1}{4}\sum_{i=0}^{i=3}\sigma_{i}$. We find the exact quantile location $\{q_{0}^{\prime},q_{1}^{\prime},q_{2}^{\prime},q_{3}^{\prime}\}$ of $\mathcal{N}(q_{\mathrm{{avg}}},\sigma^{2}_{\mathrm{{avg}}})$, then measure how much two types of quantiles are different with mean squared error.
@@ 4.3 Variance of Normal Distribution
Variance network ($\sigma^{2}_{\psi}$), introduced by Nix & Weigend (1994) and named by Mai et al. (2022), are designed to capture both the mean and Gaussian error (variance). In line with the normal assumption, the variance network allows us to estimate the variance of a Gaussian distribution. We leverage a variance network to estimate a variance for each state when computing quantile bars.
@@
$\mathcal{L}_{\sigma^{2}}(\psi)=\frac{1}{T}\sum_{i=1}^{T}\frac{1}{2\sigma^{2}_{\psi}(s_{i})}||q_{\mathrm{{avg}}}(s_{i})-G(s_{i})||^{2}+\frac{1}{2}\rm{ln}\sigma^{2}_{\psi}(s_{i})$ (6)
where $q_{\mathrm{{avg}}}(s)$ is the mean of $V^{D,\pi}_{\theta}(s)$ and $G(s)$ is the sum of sampled rewards from the state $s$ (return). This separation of prediction eliminates the need for additional hyperparameter tuning that would have been required for the simultaneous prediction of mean and variance, as observed in Mai et al. (2022). Although the numerator of Equation 6 is separated from $\psi$, the update is performed in the direction of increasing $\sigma^{2}_{\psi}$ when the error of the numerator is significant. Conversely, if the error is sufficiently small, $\sigma^{2}_{\psi}$ is decreased. $\sigma^{2}_{\psi}$ is trained to predict variance in a manner consistent with the original variance network.
@@ 4.4 Target for Distributional Value Function
We provide the quantile bars of an approximated normal distribution to $V^{D,\pi}_{\theta}$ with the variance from $\sigma^{2}_{\psi}$ above and returns. To compute the target quantile bars, we need the mean of the approximated normal distribution. The mean can be TD target or the sum of rewards $G(s)$ from $s$ to end. Here, we use return $G$ for the sake of notation simplicity. To obtain a quantile target for $V^{D,\pi}$, we first precompute the $Z$ values of the standard normal distribution. For example, when $N=4$, $Z=\{-0.841,-0.253,0.253,0.841\}$, i.e., $P(X<=Z[i])\approx 0.2*(i+1)=\frac{i+1}{N+1}$ for $i\in\{0,1,2,3\}$. Note that we use $N=8$ quantile bars for all experiments in practice, as we empirically find that $N\geq 8$ provides satisfactory performance. Then, we compute the target quantile bars $q^{\prime}_{t}$ of $V^{D,\pi}_{\theta}(s_{t})$ such that $q^{\prime}_{t}=\{R_{t}+\sigma_{\psi}Z[i]\,|\,i\in\{0,1,...,N-1\}\}$ and fit $V^{D,\pi}_{\theta}$ to $q^{\prime}_{t}$ by quantile Huber loss (Aravkin et al., 2014). The quantile Huber loss makes $V^{D,\pi}_{\theta}$ output quantile bars that match the definition of quantiles. The Huber loss (Huber, 1964) is as follows:
@@ 4.5 Uncertainty Weight
Figure 3 illustrates this when the number of quantile bars is $N=4$ for simplicity. In practice, we use $N=8$ quantile bars for all experiments, as we empirically find that $N\geq 8$ provides similar satisfactory performance. If the prediction of $V^{D,\pi}$ is reliable, it should be symmetrical with respect to the mean of the quantile bars from our normality assumption, and the location of the quantile bars should follow a normal distribution. We first find the mean $q_{\mathrm{{avg}}}(s)$ = $\frac{1}{N}\sum_{i=0}^{i=N-1}q_{i}(s)$ of the quantile bars and calculate the standard deviations $\{\sigma_{0},\sigma_{1},...,\sigma_{N-1}\}$ for each quantile such that $P(Z\leq\frac{q_{i}-q_{\mathrm{{avg}}}}{\sigma_{i}})=\frac{i+1}{N+1}$ for $i\in\{0,1,...,N-1\}$. We consider $V^{D,\pi}$ is trying to approximate $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$ where $\sigma_{\mathrm{{avg}}}=\frac{1}{N}\sum_{i=0}^{i=N-1}\sigma_{i}$.
To measure the uncertainty $w$ for policy updates, we find the locations of each quantile $\{q_{0}^{\prime}(s),...,q_{N-1}^{\prime}(s)\}$ based on $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$. The better $V^{D,\pi}$ resembles the normal distribution, the smaller the difference between these two types of quantile bars. Therefore, we use the mean sqaured error $E=\sum_{i=0}^{i=N-1}(q_{i}(s)-q_{i}^{\prime}(s))^{2}$, and we compute the uncertainty weight $w(s)$ as follows:
$w(s)=\sigma(-E*T)+0.5$ (9)
where $T>0$ is a temperature and $\sigma$ is the sigmoid function. Although Lee et al. (2020) suggest values for $T$, it seems that finding the reasonable value for $T$ still requires some tuning. We instead set a target uncertainty weight and perform a parametric search by adjusting $T$ to achieve that target weight (Appendix B).
@@
1:Input: policy $\pi$, distributional value function $V^{D,\pi}_{\theta}$, variance network $\sigma^{2}_{\psi}$, rollout buffer $B$
2:Initialize $\pi$, $V^{D,\pi}$, and $\sigma^{2}_{\psi}$
3:for $i=1$ to epoch_num do
4:     for $j=1$ to rollout_num do
5:         $a_{t}\sim\pi(\cdot|s_{t})$
6:         $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$
7:         Compute $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$ with $V^{D,\pi}(s_{t})=\{q_{0}(s_{t}),q_{1}(s_{t}),...,q_{N-1}(s_{t})\}$
8:         Find $q^{\prime}(s_{t})$ from $\mathcal{N}(q_{\mathrm{{avg}}}(s),\sigma^{2}_{\mathrm{{avg}}})$
9:         Compute the mean squared error $E=\sum_{i=0}^{N-1}(q_{i}(s_{t})-q_{i}^{\prime}(s_{t}))^{2}$
10:         Store ($s_{t},a_{t},r(s_{t},a_{t}),E,V^{D,\pi}(s_{t}),\sigma^{2}_{\psi}(s_{t})$) in $B$
11:         if $s_{t+1}$ is terminal then
12:              Reset env
13:         end if
14:     end for
15:     Perform the parametric search to find the temperature $T$ (Appendix B)
16:     Compute target quantiles for $V^{D,\pi}$ with the stored return and $\sigma^{2}_{\psi}$
17:     Minimize $\mathcal{L}_{\sigma^{2}}(\psi)$ and $\mathcal{L}_{q}(\theta)$
18:     Optimize $\pi$ with a policy objective scaled by $w$
19:end for
4.6 Policy Update with Uncertainty Weight
We leverage the uncertainty weights $w$ to update the policy to mainly focus on values for which the current $V^{D,\pi}$ predicts correctly. Each reinforcement learning algorithm has a policy objective function. We scale the objective function with the uncertainty weight to adjust the influence of the gradient for each sample. For TRPO (Schulman et al., 2015a),
[…]
Table 1: Ablation study of TRPO and PPO. Each entry represents an average return and a standard error, derived from 30 different settings, with each setting running for 100 episodes. $V^{D,\pi}$ (QR) represents the distributional value function trained with the Huber quantile regression. This table illustrates an improvement trend across the scores with the addition of each component from the baseline to MC-CLT. Also, MC-CLT results generally better than the Ensemble method, while using fewer weights and reduced training time. An asterisk (*) indicates a statistically significant improvement between MC-CLT and Ensemble method at a significance level of $0.05$
@@ 4.6 Policy Update with Uncertainty Weight
Swimmer Walker2d InvertedDoublePendulum LunarLander(C)
@@
Swimmer Walker2d InvertedDoublePendulum LunarLander(C)
@@ 5.1 Experiment Setups
Our implementation is based on Spinning Up (Achiam, 2018), an open-source resource by OpenAI that provides implementations of several reinforcement learning algorithms and tools to help researchers get started with deep RL. For our experiments, we chose two representative deep reinforcement learning algorithms, Policy Optimization (PPO) (Schulman et al., 2017) and Trust Region Policy Optimization (TRPO) (Schulman et al., 2015a), to evaluate our method. The advantage is computed by GAE (Schulman et al., 2015b).
We use the default hyperparameters such as learning rate and batch size. All policies have a two-layer tanh network with 64 x 32 units, and all of the value function and the distributional value function has a two-layer ReLU network with 64 x 64 units or 128 x 128 units for all environments. All networks are updated with Adam optimizer (Kingma & Ba, 2014). We evaluate our method on continuous OpenAI gym Box2D (Brockman et al., 2016) and MuJoCo tasks (Todorov et al., 2012), as these environments have continuous action spaces and dense reward functions to use the normal approximation of MC-CLT (Theorem 1).
@@ 5.2 Detailed Ablation Analysis and Comparison
Table 1 represents the experimental results across 8 environments. The first section of the table provides results from policies trained using TRPO, while the second section gives results from policies trained with PPO. As we incrementally add each component from the baseline to MC-CLT, a consistent trend of performance improvement is observed. Also, our method shows better results compared to the ensemble-based approach in general. In 10 out of the 16 settings, we achieve statistically significant improvements between MC-CLT and the Ensemble method, as indicated by an asterisk (*) in the table, at a significance level of $0.05$. It’s worth noting that the ensemble-based approach with TRPO exhibits better mean performances in three settings: HalfCheetah, Ant, and InvertedDoublePendulum. However, the differences are not statistically significant.
@@
In addition, MC-CLT utilizes fewer weights and achieves faster training time, as shown in (Table 2). Table 2 presents a relative comparison with the other two methods based on the baseline. Specifically, when the training time for PPO is set to 1, the MC-CLT method requires 1.5 to 1.9 times more training time compared to the baseline. Similarly, the ensemble-based approach takes 2.8 to 3.1 times more time. Additionally, the MC-CLT method requires approximately 2.0 to 2.2 times more weights, while the ensemble-based approach uses 5 times as many parameters. We use 8 quantile bars for all experiments, as we empirically find that using 8 or more quantile bars provides smiliar satisfactory performance.
While MC-CLT requires hyperparameters for the number of quantile bars and the variance network, it necessitates fewer hyperparameters compared to previous methods (Lee et al., 2020; Mai et al., 2022). We observe satisfactory performance when the number of quantiles is set to $\geq 8$ and the size of the variance network equals to that of the distributional value function. In practice, we fix the number of quantiles at 8 for all experiments. Consequently, aside from the aspects controlling uncertainty, MC-CLT shares the same hyperparameters as the baseline method. The detailed hyperparameter settings are discussed in Appendix A.
5.3 Seeking Accuracy vs. Seeking Exploration
MC-CLT and the ensemble-based methods seek accuracy by reducing weights for samples with high uncertainty. This raises exploration concerns regarding states where the current distributional value function fails to predict accurately. If such states are not given sufficient weight, the policy may never cover them effectively. To address this concern, we conduct additional experiments where states with higher uncertainty receive higher weights to encourage the agent to explore these states more. We compute the weight with $w(s)=E*T+0.5$, where $E=\sum_{i=0}^{N-1}(q_{i}(s)-q_{i}^{\prime}(s))^{2}$, and run Algorithm 2 as in MC-CLT to find the temperature $T$ (where left and right are adjusted in the opposite direction from accuracy seeking). Table 3 shows the comparisons between accuracy-seeking and exploration-seeking approaches. In all experiments, the accuracy-seeking approach shows better performance. This is likely because the environments use highly engineered reward functions, providing sufficient feedback without requiring additional exploration. For tasks where discovering novel states is important, such as in Atari games (Bellemare et al., 2013), the exploration-seeking approach would have a positive effect on performance (Obando-Ceron et al., 2023; Schaul et al., 2022).